OpenStudy (anonymous):

can someone explain to me step by step how to do this question: three pumps can transfer 4150 gal of water per day when working at the same time. Pumps A and B together can transfer 3200 gal per day. Pumps A and C together can transfer 2900 gal per day. How many gallons can each pump transfer working alone? Translate the equation and solve.

OpenStudy (phi):

how much do A and B transfer in a day. Can you write that as an equation?

OpenStudy (anonymous):

x= pump A y= pump B z= pump C x+y+z= 4150 x+y= 3200 x+z= 2900 i dont know what to do from here.

OpenStudy (phi):

that is a good start. (Though I would have used A, B, C as the variables, why add to the confusion?)

OpenStudy (phi):

We have 3 equations and 3 unknowns I would use elimination. write it as 1x+1y+1z= 4150 1x+1y+0z= 3200 1x+0y+1z= 2900 if we ignore for the moment the x,y,z 1 1 1 4150 1 1 0 3200 1 0 1 2900 Elimination means make the first column 1 0 0 now multiply the 1st row by -1 and add the result to the 2nd row Can you do that?

OpenStudy (anonymous):

z=950

OpenStudy (phi):

1 1 1 4150 (mult by -1 to get -1 -1 -1 -4150.) 1 1 0 3200 add -1 -1 -1 -4150 to get 0 0 -1 -950 1 0 1 2900 the new problem is 1 1 1 4150 0 0 -1 -950 1 0 1 2900 to get rid of the 1 in the last row, do the same thing. mult the first row by -1 and add to the last row 1 1 1 4150 0 0 -1 -950 1 0 1 2900 add -1 -1 -1 -4150 to get 0 -1 0 -1250 we now have 1 1 1 4150 0 0 -1 -950 0 -1 0 -1250 if we add back the x,y,z this means x+y+z= 4150 -z= -950 -y= -1250

OpenStudy (anonymous):

x=1950

OpenStudy (anonymous):

thank you. can you help me with another question?

OpenStudy (phi):

Looks like you got the answers, Post it.

OpenStudy (anonymous):

The members of the graduating class at Grant High School arranged themselves in 8 rows with the same number of students in each row. When they rearranged themelves into 9 rows, they were 3 fewer members in each row than were before. How many members are there in the class?

OpenStudy (phi):

make two equations. x is the number of students in the 8 rows x-3 is the number when they use 9 rows y is total number of students

OpenStudy (phi):

let me be clearer. x is the number of students in 1 row, when they assembled into 8 rows so 8x= y

OpenStudy (anonymous):

would the eqations be x-3=9 and 8x-y=0?

OpenStudy (phi):

I was thinking 8x=y 9(x-3)= y

OpenStudy (anonymous):

i dont understand. the answers i can choose from are a)144 b)192 c)216 d)243 e)388

OpenStudy (anonymous):

e)288*

OpenStudy (phi):

if you have x people in 8 rows then you have 8x people they then tell us When they rearranged themelves into 9 rows, they were 3 fewer members in each row so if there were x people in each row, now there are x-3 people in each row the total number of people is 9(x-3) if we call y the total number we have 8x=y 9(x-3)=y

OpenStudy (phi):